OpenStudy (anonymous):

lim 3+2x+4x^2/x^2 x->inf I know as x approaches inf the x^2 will increase much faster making the lower degree terms insignificant. thats where I am stuck

OpenStudy (anonymous):

End behavior model of the numerator:\[x^2\]End behavior model of the denominator:\[x^2\]Restatement of the limit approaching infinity:\[\lim_{x\rightarrow\infty}\dfrac{x^2}{x^2}=\lim_{x\rightarrow\infty}1=1\]

OpenStudy (anonymous):

I thought as much its just that I wasn't sure if you could cancel the variables like that thanks

OpenStudy (anonymous):

You can because with extremely large values of x, the numerator and the denominator are essentially equal, and thus the quotient will be basically 1.

OpenStudy (anonymous):

what about the 4 in the 4x^2

OpenStudy (anonymous):

Ach! My apologies! I missed seeing that. You are right! We thus have \[\lim_{x\rightarrow\infty}\dfrac{4x^2}{x^2}=4\]Good spotting! Do watch out for those! :)

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

I got another question I will post it so you can get another medal